# Split Matching for Inductive Zero-shot Semantic Segmentation

### v2

@@ Abstract
Zero-shot Semantic Segmentation (ZSS) aims to segment categories that are not annotated during training. While fine-tuning vision-language models has achieved promising results, these models often overfit to seen categories due to the lack of supervision for unseen classes. As an alternative to fully supervised approaches, query-based segmentation has shown great latent in ZSS, as it enables object localization without relying on explicit labels. However, conventional Hungarian matching, a core component in query-based frameworks, needs full supervision and often misclassifies unseen categories as background in the setting of ZSS. To address this issue, we propose Split Matching (SM), a novel assignment strategy that decouples Hungarian matching into two components: one for seen classes in annotated regions and another for latent classes in unannotated regions (referred to as unseen candidates). Specifically, we partition the queries into seen and candidate groups, enabling each to be optimized independently according to its available supervision. To discover unseen candidates, we cluster CLIP dense features to generate pseudo masks and extract region-level embeddings using CLS tokens. Matching is then conducted separately for the two groups based on both class-level similarity and mask-level consistency. Additionally, we introduce a Multi-scale Feature Enhancement (MFE) module that refines decoder features through residual multi-scale aggregation, improving the model’s ability to capture spatial details across resolutions. SM is the first to introduce decoupled Hungarian matching under the inductive ZSS setting, and achieves state-of-the-art performance on two standard benchmarks.
@@ 1 Introduction
Semantic segmentation [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2024b)Chen, Deguchi, Zhang, Zheng, and Murase] serves as a fundamental task for fine-grained scene understanding in various applications, such as autonomous driving, robotic perception, and medical imaging [Zheng et al.(2022)Zheng, Fu, Xie, Chen, Wang, and Sham, Chen et al.(2022)Chen, Fu, Xie, Zheng, Geng, and Sham, Zheng et al.(2024a)Zheng, Luo, Fu, Liu, and Wang, Xie et al.(2023)Xie, Fu, Zheng, Zheng, Sham, and Wang]. Existing approaches can be broadly classified into feature-based and query-based methods. Feature-based methods treat semantic segmentation as a per-pixel classification problem, where dense features extracted from the backbone are directly finetuned for pixel-wise label prediction. In contrast, query-based segmentors employ a set of discrete, learnable vectors, referred to as queries, to jointly predict class labels and class-agnostic masks. These queries are passed through a transformer decoder to produce class scores and simultaneously interact with dense backbone features to generate the corresponding masks.
@@
[…]
To address the limitations of existing approaches, particularly their tendency to overfit to seen classes, we explore query-based segmentation as a more effective alternative for zero-shot settings. Unlike feature-based methods that classify each pixel independently, query-based methods treat each object as a learnable query and perform mask-level classification, enabling better modeling of object-level semantics and localization of latent classes without requiring explicit labels. Despite this advantage, current query-based methods struggle in ZSS scenarios, where the absence of annotations for unseen classes during training causes queries to be predominantly assigned to seen categories. As a result, queries corresponding to unseen objects remain unmatched, are treated as background, and thus fail to detect unseen categories. As illustrated in the left of Fig. 1, this matching imbalance introduces a strong bias toward seen classes, leaving the model unable to train or preserve representations for unseen classes, limiting the full latent of query-based segmentation for ZSS.
To address the challenge of segmenting unseen classes, we propose Split Matching (SM) for query-based segmentation. Specifically, we divide the queries into two groups: seen queries, which segment annotated seen classes, and candidate queries, which target latent classes in unannotated regions (referred to as unseen candidates). We first apply multi-scale K-Means clustering [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase] on CLIP dense features to generate pseudo masks that localize unseen candidates. The corresponding image patches are then cropped and fed into the frozen CLIP visual encoder to obtain CLS tokens, which are subsequently fused with semantic embeddings to form joint vision-language representations. We compute class similarity between the joint embeddings and query predictions, and measure mask similarity using the pseudo masks. Hungarian matching is then applied separately to the seen and candidate queries based on the combined similarities. To further enhance feature quality, we introduce a Multi-scale Feature Enhancement (MFE) module, which refines the transformer decoder’s key and value via residual multi-scale aggregation.
Different from existing methods that rely on dense features extracted from CLIP [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase] and struggle to optimize unannotated regions due to the absence of explicit supervision, our method introduces Split Matching (SM), which separates queries into seen and candidate groups, enabling targeted label assignment even in unannotated areas. This design effectively mitigates the common issue of misclassifying unseen objects as background—a key limitation of prior methods. Moreover, unlike open-vocabulary segmentation approaches [Yu et al.(2023)Yu, He, Deng, Shen, and Chen, Xu et al.(2023)Xu, Zhang, Wei, Hu, and Bai] that perform Hungarian matching with fully annotated data, our method operates entirely under the zero-shot setting, without requiring any pixel-level labels for unseen classes. To the best of our knowledge, we are the first to explicitly separate seen and candidate queries via Hungarian matching under the inductive ZSS. To summarize, our contributions are:
- We propose Split Matching (SM), the first query-based assignment framework for zero-shot segmentation that separates seen and unseen candidates using pseudo masks derived from CLIP features without requiring fully labeled datasets.
- We introduce a Multi-scale Feature Enhancement (MFE) module that enriches decoder features via residual fusion, improving query interaction and segmentation quality.
- Our method achieves state-of-the-art performance on PASCAL VOC and COCO-Stuff, demonstrating strong generalization to unseen classes and balanced performance.
@@ 2 Related Works
Semantic Segmentation assigns a class label to each pixel in an image [Zheng et al.(2023b)Zheng, Zhu, Liu, Cao, Fu, and Wang, Zheng et al.(2023a)Zheng, Pan, Luo, and Wang, Zheng et al.(2024e)Zheng, Zhou, Vasilakos, and Wang, Zheng et al.(2024b)Zheng, Lyu, and Wang, Zheng et al.(2024d)Zheng, Zhou, Vasilakos, and Wang, Zheng et al.(2024c)Zheng, Lyu, Zhou, and Wang, Zhu et al.(2024)Zhu, Xiao, Shi, Xu, and Zheng, Liao et al.(2025b)Liao, Zheng, Lyu, Xue, Cao, Wang, Yang, and Hu, Liao et al.(2025a)Liao, Lei, Zheng, Moon, Wang, Wang, Paudel, Gool, and Hu, Zhong et al.(2025)Zhong, Zheng, Liao, Lyu, Chen, Wu, Zhang, and Hu]. Traditional CNN-based methods [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] enhance per-pixel classification through dilated convolutions and multi-scale context aggregation but struggle to capture long-range dependencies. With the rise of Vision Transformers [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.], encoder-based models [Xie et al.(2021)Xie, Wang, Yu, Anandkumar, Alvarez, and Luo, Guo et al.(2022)Guo, Lu, Hou, Liu, Cheng, and Hu] have shown strong capabilities in modeling global context. Inspired by DETR [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko], query-based frameworks such as MaskFormer [Cheng et al.(2021a)Cheng, Schwing, and Kirillov] and Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] reformulate segmentation as set prediction using object queries and Hungarian matching. Although these models achieve impressive results under full supervision, they are not directly applicable to zero-shot segmentation (ZSS) due to their reliance on complete annotations and inherent bias toward seen classes. To address this, we propose Split Matching, the first method to explicitly separate seen and candidate queries under the inductive ZSS setting. By decoupling the matching process for annotated and unannotated regions, our approach enables targeted supervision and significantly improves generalization to unseen categories, even in the absence of pixel-level annotations.
@@
Zero-shot and Open-Vocabulary Segmentation. Although semantic segmentation has made remarkable progress, it still relies heavily on large-scale pixel-level annotations, which are expensive and time-consuming to obtain. To alleviate this, two related directions have emerged: zero-shot segmentation (ZSS) [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Han et al.(2023)Han, Zhong, Li, Han, and Ma, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase, Chen et al.(2024a)Chen, Deguchi, Zhang, and Murase, Ye et al.(2024)Ye, Oh, et al.] and open-vocabulary segmentation (OVS) [Xu et al.(2023)Xu, Zhang, Wei, Hu, and Bai, Yu et al.(2023)Yu, He, Deng, Shen, and Chen, Wu et al.(2024)Wu, Zhang, Xu, Jin, Li, Liu, and Loy, Zhou et al.(2022)Zhou, Loy, and Dai]. Both paradigms aim to improve generalization by leveraging large-scale vision-language models such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], either by introducing lightweight adapters [Houlsby et al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De Laroussilhe, Gesmundo, Attariyan, and Gelly] or designing task-specific visual prompts [Jia et al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim]. Despite this shared goal, their setups differ fundamentally. ZSS is defined under a partially labeled training regime, where only a subset of classes is annotated and the rest are marked as ignored. Evaluation is conducted on the same domain, including both seen and unseen categories. OVS, on the other hand, assumes access to fully labeled training data and evaluates on datasets with novel categories and distribution shifts. While OVS benefits from full supervision and can leverage existing segmentation architectures, ZSS faces the challenge of incomplete supervision. Unannotated regions corresponding to unseen classes are often treated as background during training, making it difficult for models to learn discriminative representations for unseen concepts. To tackle this issue, we propose Split Matching (SM), a label assignment strategy tailored for query-based models in the zero-shot setting. SM dynamically aligns predicted queries with pseudo labels derived from external vision-language features, enabling the discovery of unseen objects even in the absence of ground-truth annotations.
@@
Figure 2: Overview of the proposed method. During training, input images are fed into a trainable backbone to extract dense features, while a frozen CLIP encoder provides CLIP features. These are used to generate class embeddings $\textbf{C}_{u}$ and pseudo masks $\textbf{Y}_{p}$ for both seen and unseen classes. The dense features serve as keys and values in a transformer decoder, which interacts with queries Q to predict masks and query features. These are aligned with pseudo labels via the Split Matching. A Multi-scale Feature Enhancement module further refines features with auxiliary loss $\mathcal{L}_{MFE}$.
@@ 3.1 Preliminaries
Task Definition. ZSS aims to segment both seen and unseen classes without pixel-level annotations for unseen classes after training. Formally, let $\mathcal{D}=\{\textbf{X}^{i},\textbf{Y}_{s}^{i}\}_{i=1}^{M}$ represent a dataset of images X and their pixel-level annotations $\textbf{Y}_{s}$ for seen classes. Additionally, let $\textbf{A}\in\mathbb{R}^{N\times C}$ denote the semantic (text) embeddings of all classes, divided into seen $\textbf{A}_{s}\in\mathbb{R}^{N_{s}\times C}$ and unseen $\textbf{A}_{u}\in\mathbb{R}^{N_{u}\times C}$, such that $(\textbf{A}_{s}\cap\textbf{A}_{u}=\varnothing)$ and $N_{s}+N_{u}=N$ where $C$ indicates the channel number and $N$ indicates the number of classes in the dataset. Our method applies the Inductive settings, where unseen embeddings $\textbf{A}_{u}$ are inaccessible during training and evaluates model performance on seen and unseen classes during inference. Meanwhile, all training images are preserved during training, and regions corresponding to unseen classes are consistently labeled as “ignored”, ensuring that no unseen information is used.
Method Overview. Our core idea is to mitigate the seen-class bias caused by incomplete annotations by decoupling the optimization of seen and unannotated regions, thereby enabling better discovery of unseen classes without degrading seen-class performance. As illustrated in Fig. 2, we divide the query space into two parts: seen queries and candidate queries, which are responsible for segmenting seen classes in the annotated areas and unseen candidates in the unannotated areas, respectively. Given an input image, a trainable backbone extracts dense features that serve as keys and values for a transformer decoder. A set of randomly initialized queries $\mathbf{Q}$ interacts with these features to produce segmentation outputs, which are then split into $\mathbf{Q}_{s}$ (seen queries) and $\mathbf{Q}_{u}$ (candidate queries). To guide $\mathbf{Q}_{u}$ towards discovering unseen candidates, the input image is also passed through a frozen CLIP encoder. From the CLIP features, we generate pseudo masks $\mathbf{Y}_{u}$ and corresponding class embeddings $\mathbf{C}_{u}$ by extracting the CLS tokens of the unseen candidates. We then introduce Split Matching (Sec.3.2), which assigns labels to both seen and candidate queries by computing similarity between model predictions, pseudo masks, and class embeddings. To further enhance feature quality, we propose a Multi-scale Feature Enhancement module (Sec.3.3), which refines the dense features extracted by the backbone.
Although our method utilizes features from unannotated regions, it strictly adheres to the inductive zero-shot setting. At the beginning of training, all unannotated regions are uniformly treated as ‘ignored’ without introducing any class-specific supervision or bias. Unseen candidates are identified in a fully self-supervised manner, without any assumptions regarding the number, identity, or distribution of unseen classes. As a result, the model remains completely agnostic to unseen classes throughout the entire training process, ensuring that no unseen-related information is leaked.

### v3

@@ Abstract
Zero-shot Semantic Segmentation (ZSS) targets the segmentation of unseen classes, i.e., classes not annotated during training. While fine-tuned vision-language models show promise, they often overfit to seen classes due to the lack of supervision. Query-based methods offer strong potential by enabling object localization without explicit labels, but conventional approaches assume full supervision and thus tend to misclassify unseen classes as background in ZSS settings. To address this issue, we propose Split Matching (SM), a novel assignment strategy that decouples Hungarian matching into two components: one for seen classes in annotated regions and another for latent classes in unannotated regions (referred to as unseen candidates). Specifically, we split the queries into seen and candidate queries, enabling each to be optimized independently according to its available supervision. To discover unseen candidates, we cluster CLIP dense features to generate pseudo masks and extract region-level embeddings using CLS tokens. Matching is then conducted separately for the two groups based on both class and mask similarity. Additionally, we introduce a Multi-scale Feature Enhancement (MFE) module that refines decoder features through residual multi-scale aggregation, improving the model’s ability to capture spatial details across resolutions. Besides, we also introduce a Random Query (RQ) that injects a few randomly initialized queries during inference. These queries act as unbiased probes, enriching the density and diversity of the query space. Our method is the first to introduce decoupled Hungarian matching under the inductive ZSS setting, and achieves 0.8% and 1.1% higher hIoU on two ZSS benchmarks.
@@ 1 Introduction
Semantic segmentation [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2024b)Chen, Deguchi, Zhang, Zheng, and Murase, Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Cheng et al.(2021a)Cheng, Schwing, and Kirillov, Zheng et al.(2025)Zheng, Luo, Zhou, and Wang] serves as a fundamental task for computer vision. Existing approaches can be broadly classified into feature-based and query-based methods. Feature-based [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2024b)Chen, Deguchi, Zhang, Zheng, and Murase] methods treat semantic segmentation as a per-pixel classification problem, where dense features extracted from the backbone are directly finetuned for pixel-wise label prediction. In contrast, query-based segmentors [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Cheng et al.(2021a)Cheng, Schwing, and Kirillov] employ a set of discrete, learnable vectors, referred to as queries, to jointly predict class labels and class-agnostic masks. These queries are passed through a transformer decoder to produce class scores and interact with dense backbone features to generate the corresponding masks.
However, achieving high performance in semantic segmentation typically requires large-scale datasets with pixel-level annotations [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele], which are costly to obtain. To reduce annotation demands, Zero-shot Semantic Segmentation (ZSS) [Bucher et al.(2019)Bucher, Vu, Cord, and Pérez, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase, Chen et al.(2024a)Chen, Deguchi, Zhang, and Murase] has emerged, aiming to segment unseen classes which are not annotated during training but must be segmented at test time, by transferring knowledge from seen classes, i.e., classes with available training annotations. Recent advances in Vision-Language Models (VLMs) [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Zhang et al.(2023)Zhang, Huang, Ma, Li, Luo, Xie, Qin, Luo, Li, Liu, et al.] drives ZSS by enabling the transfer of vision-language alignment to segmentation tasks [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Han et al.(2023)Han, Zhong, Li, Han, and Ma]. Existing methods typically adapt CLIP via adapters [Houlsby et al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De Laroussilhe, Gesmundo, Attariyan, and Gelly] or prompts [Jia et al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim] to handle seen classes, while relying on CLIP’s generalization ability for unseen ones. However, such strategies typically rely on CLIP’s zero-shot capability for unseen classes without additional training signals, which leads to overfitting to seen classes and poor generalization on unseen classes.
To mitigate overfitting to seen classes, we explore query-based segmentation as a more effective solution. Unlike feature-based methods that classify pixels independently, query-based approaches treat each object as a learnable query and perform mask-level classification, enabling better object-level reasoning and localization without explicit supervision. However, despite their ability to localize potential objects, existing query-based methods struggle in ZSS. Without annotations for unseen classes during training, queries for novel objects are often misclassified as seen classes or background. As shown on the left of Fig. 1, the imbalance biases the model to seen classes and prevents it from learning useful representations for unseen ones, limiting the full potential of query-based segmentation in ZSS.
To improve the ability of query-based methods to segment unseen classes, we propose a method called Split Matching (SM). Specifically, we divide the queries into two groups: seen queries, which segment annotated seen classes, and candidate queries, which target latent classes in unannotated regions (referred to as unseen candidates). We first apply multi-scale K-Means clustering [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase] on CLIP dense features to generate pseudo masks that localize unseen candidates. The corresponding image patches are then cropped and fed into the frozen CLIP visual encoder to obtain CLS tokens, which are subsequently fused with semantic embeddings to form joint class embedding. We compute class similarity between the joint embeddings and query predictions, and measure mask similarity using the pseudo masks. Hungarian matching is then applied separately to the seen and candidate queries based on the combined similarities. Although SM facilitates the adaptation of query-based approaches to zero-shot segmentation, the key and value features in transformer decoders, responsible for associating queries with relevant image regions, remain suboptimal. These features often lack sufficient spatial granularity and discriminative power, causing noisy query-to-region associations and reduced matching accuracy. Therefore, we introduce a Multi-scale Feature Enhancement (MFE). It enhances the visual feature with multi-scale context and applies spatial normalization, refining the transformer decoder’s key and value through residual multi-scale aggregation. Moreover, we propose a Random Query (RQ) strategy that introduces a small number of randomly initialized queries at inference time. Unlike trained queries, these random queries serve as unbiased exploratory probes that complement the existing query set. By diversifying the query space and increasing its sampling density, RQ reduces blind spots and encourages the decoder to discover additional object regions that would otherwise be missed.
@@
Different from existing methods that rely on dense features extracted from CLIP [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase] and struggle to optimize unannotated regions due to the absence of explicit supervision, our method introduces Split Matching (SM), which separates queries into seen and candidate groups, enabling targeted label assignment even in unannotated areas. This design effectively mitigates the common issue of misclassifying unseen objects as background, a key limitation of prior methods. Moreover, unlike open-vocabulary segmentation approaches [Yu et al.(2023)Yu, He, Deng, Shen, and Chen, Xu et al.(2023)Xu, Zhang, Wei, Hu, and Bai] that perform Hungarian matching with fully annotated data, our method operates entirely under the zero-shot setting, without requiring any pixel-level labels for unseen classes. To the best of our knowledge, we are the first to explicitly separate seen and candidate queries via Hungarian matching under the inductive ZSS. To summarize, our contributions are:
• We propose Split Matching (SM), a novel query-based assignment framework for zero-shot segmentation. SM explicitly separates seen and unseen queries and matches them independently using pseudo masks derived from CLIP dense features.
• We introduce a Multi-scale Feature Enhancement (MFE) module, which enriches decoder features via residual multi-scale fusion. Additionally, we propose a Random Query (RQ) strategy that increases query density at inference time to uncover more latent objects.
• Our method achieves 0.8% and 1.1% higher hIoU on PASCAL VOC and COCO-Stuff benchmarks under the zero-shot setting.
@@ 2 Related Works
Semantic Segmentation assigns a class label to each pixel in an image. Traditional CNN-based methods [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] enhance per-pixel classification through dilated convolutions and multi-scale context aggregation but struggle to capture long-range dependencies. With the rise of Vision Transformers [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.], encoder-based models [Xie et al.(2021)Xie, Wang, Yu, Anandkumar, Alvarez, and Luo, Guo et al.(2022)Guo, Lu, Hou, Liu, Cheng, and Hu] have shown strong capabilities in modeling global context. Inspired by DETR [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko], query-based frameworks such as MaskFormer [Cheng et al.(2021a)Cheng, Schwing, and Kirillov] and Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] reformulate segmentation as set prediction using object queries and Hungarian matching. Although these models achieve impressive results under full supervision, they are not directly applicable to zero-shot segmentation (ZSS) due to their reliance on complete annotations and inherent bias toward seen classes. To address this, we propose Split Matching, the first method to explicitly separate seen and candidate queries under the inductive ZSS setting. By decoupling the matching process for annotated and unannotated regions, our approach enables targeted supervision and significantly improves generalization to unseen classes, even in the absence of pixel-level annotations.
@@
Zero-shot and Open-Vocabulary Segmentation. Although semantic segmentation has made remarkable progress, it still relies heavily on large-scale pixel-level annotations, which are expensive and time-consuming to obtain. To alleviate this, two related directions have emerged: zero-shot segmentation (ZSS) [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Han et al.(2023)Han, Zhong, Li, Han, and Ma, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase, Chen et al.(2024a)Chen, Deguchi, Zhang, and Murase, Ye et al.(2024)Ye, Oh, et al.] and open-vocabulary segmentation (OVS) [Xu et al.(2023)Xu, Zhang, Wei, Hu, and Bai, Yu et al.(2023)Yu, He, Deng, Shen, and Chen, Wu et al.(2024)Wu, Zhang, Xu, Jin, Li, Liu, and Loy, Zhou et al.(2022)Zhou, Loy, and Dai]. Both paradigms aim to improve generalization by leveraging large-scale vision-language models such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], either by introducing lightweight adapters [Houlsby et al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De Laroussilhe, Gesmundo, Attariyan, and Gelly, Quan et al.(2025)Quan, Chen, Deguchi, Sun, Zhang, Li, and Murase] or designing task-specific visual prompts [Jia et al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim]. Despite this shared goal, their setups differ fundamentally. ZSS is defined under a partially labeled training regime, where only a subset of classes is annotated and the rest are marked as ignored. Evaluation is conducted on the same domain, including both seen and unseen classes. OVS, on the other hand, assumes access to fully labeled training data and evaluates on datasets with novel classes and distribution shifts. While OVS benefits from full supervision and can leverage existing segmentation architectures, ZSS faces the challenge of incomplete supervision. Unannotated regions corresponding to unseen classes are often treated as background during training, making it difficult for models to learn discriminative representations for unseen concepts. To tackle this issue, we propose Split Matching (SM), a label assignment strategy tailored for query-based models in the zero-shot setting. SM dynamically aligns predicted queries with pseudo labels derived from external vision-language features, enabling the discovery of unseen objects even in the absence of ground-truth annotations.
@@
Figure 2: Training pipeline overview of the proposed method. During training, input images are fed into a trainable backbone to extract dense features F, while a frozen CLIP encoder provides CLIP features. These are used to generate class embeddings $\textbf{C}_{u}$ and pseudo masks $\textbf{Y}_{p}$ for both seen and unseen classes. The dense features serve as keys and values in a transformer decoder, which interacts with queries Q to predict masks and query features. These are aligned with pseudo masks via the Split Matching. A Multi-scale Feature Enhancement module further refines F with $\mathcal{L}_{MFE}$.
@@ 3.1 Preliminaries
Task Definition. ZSS aims to segment both seen and unseen classes without annotations for unseen classes during training. Formally, let $\mathcal{D}=\{\textbf{X}^{i},\textbf{Y}_{s}^{i}\}_{i=1}^{M}$ represent a dataset of images X and their pixel-level annotations $\textbf{Y}_{s}$ for seen classes, and $M$ is the dataset size. Additionally, let $\textbf{A}\in\mathbb{R}^{N\times C}$ denote the semantic (text) embeddings of all classes, divided into seen $\textbf{A}_{s}\in\mathbb{R}^{N_{s}\times C}$ and unseen $\textbf{A}_{u}\in\mathbb{R}^{N_{u}\times C}$, such that $(\textbf{A}_{s}\cap\textbf{A}_{u}=\varnothing)$ and $N_{s}+N_{u}=N$ where $C$ indicates the channel number and $N$ is the number of classes in the dataset. Our method applies the Inductive settings, where unseen embeddings $\textbf{A}_{u}$ are inaccessible during training and evaluates model performance on seen and unseen classes during inference. Meanwhile, all training images are preserved during training, and regions corresponding to unseen classes are consistently labeled as “ignored”, ensuring that no unseen information is used. Although our method utilizes features from unannotated regions, it strictly adheres to the inductive zero-shot setting. At the beginning of training, all unannotated regions are uniformly treated as ‘ignored’ without introducing any class-specific supervision or bias. Unseen candidates are identified in a fully self-supervised manner, without any assumptions regarding the number, identity, or distribution of unseen classes. Therefore, the model remains agnostic to unseen classes during training, ensuring that no unseen-related information is leaked.
Method Overview. Our core idea is to mitigate seen-class bias from incomplete annotations by decoupling the optimization of seen and unannotated regions, allowing better discovery of unseen classes without harming seen-class performance. As shown in Fig.2, we divide the query space into seen queries $\mathbf{Q}_{s}$ and candidate queries $\mathbf{Q}_{u}$, responsible for segmenting annotated seen classes and unseen candidates, respectively. Given an input image, a trainable backbone extracts dense features, which interact with a set of randomly initialized queries through a transformer decoder. These queries are split into $\mathbf{Q}_{s}$ and $\mathbf{Q}_{u}$, with the latter guided by pseudo masks and class embeddings for unseen candidates $\mathbf{C}_{u}$ derived from a frozen CLIP encoder. We then propose Split Matching (Sec.3.2) to assign labels to both query types via similarity with model outputs, pseudo masks, and class embeddings. Additionally, a Multi-scale Feature Enhancement module (Sec. 3.3) is introduced to refine backbone features, and a Random Query (RQ) strategy (Sec. 3.4) is introduced during inference.